OpenStudy (anonymous):

I am confused about graphical transformations on linear equations. Do we still use the y=mx+b formula? Can someone show me an example?

OpenStudy (anonymous):

No. If you're trying to find a translation, or dilation, etc... you use differen't formulas. y=mx+b is finding linear equations.

OpenStudy (anonymous):

If you're finding transformations, they're differen't

OpenStudy (anonymous):

I think...

OpenStudy (anonymous):

so what is some examples of the other formulas?

OpenStudy (anonymous):

Like... for dilations, you need to figure out how to make it larger, technically, you just find what k=... and then multiply

OpenStudy (anonymous):

for example -2,7

OpenStudy (anonymous):

and if k=3

OpenStudy (anonymous):

it would be -2 time 3 and 7 time 3

OpenStudy (anonymous):

y=mx+b is basically invalid

OpenStudy (anonymous):

it all depends on whether you want to make it smaller or larger

OpenStudy (anonymous):

Yeah...

OpenStudy (anonymous):

now im more confused.. :(

OpenStudy (zepp):

m is the slope, when it is positive, the function is increasing, and when it's getting bigger, the slope becomes more and more vertical. y = mx + b b is the y-intercept, it translate the function upward or downward

OpenStudy (anonymous):

so you do still use the formula?

OpenStudy (zepp):

Let me give you some examples; Let's say we have this very basic function, y = x http://puu.sh/s8N7

OpenStudy (zepp):

If we change the slope to 2, http://puu.sh/s8Nl It would be like this.

OpenStudy (zepp):

If you want to translate the function upward, let's add something to our function, y = 2x + 5 so our y-intercept would be at 5 http://puu.sh/s8O0

OpenStudy (anonymous):

and say the slope was -2? it would go the opposite way? or would that make it a reflection?

OpenStudy (zepp):

If the slope is -2, let's compare y = 2x and y = -2x http://puu.sh/s8Ox In this case, you can say that the function is reflected on the x-axis.

OpenStudy (zepp):

When the slope is positive, the function is strictly increasing, and when the slope is negative, the function is strictly decresing. :) Have any other questions?

OpenStudy (anonymous):

oooh thanks that makes it a lot more clear. yes does linear have like stretches or shrinks?

OpenStudy (anonymous):

like some other functions do

OpenStudy (anonymous):

no you use the transformation rules:

OpenStudy (zepp):

If the slope's value is getting bigger and bigger, we can say that the function has been streched, yes.

OpenStudy (zepp):

and you don't 'really' need to learn the rules that babarica posted by heart, since this is all about observations, but it is still helpful if you need references :)

OpenStudy (anonymous):

okay thanks so w what you said about stretching if the slope gets bigger..would it be compression if the slope gets smaller?

OpenStudy (zepp):

Yes, in fact, the absolute value of the slope getting bigger = stretch If we have y = -2x and y = -5x The function has been streched, even though -5 is smaller than -2, so we have to take a absolute value of the slope to compare.

OpenStudy (anonymous):

oooh okay makes sense thanks soo much for your help Zepp! :)

OpenStudy (zepp):

Here: http://puu.sh/s8Vg Np ;)

OpenStudy (anonymous):

thanks!